# Discrete Approximate Circle Bundles

### v2

@@ 2.5 The Universal Circle Bundle And Dimensionality Reduction
$\text{Bun}_{\mathbb{S}^{1}}(B)\quad\leftrightarrow\quad\text{Prin}_{O(2)}(B)\quad\leftrightarrow\quad\text{Vect}_{2,\mathbb{R}}(B)$
for any paracompact space $B$. It follows that up to homotopy, every circle bundle $\xi$ over $B$ can be realized (up to isomorphism) as the sphere bundle of a rank-2 real vector bundle $\xi_{\text{vect}}$.
A model for the classifying space of real rank-2 vector bundles is the infinite Grassmann manifold $\text{Gr}(2)=\text{Gr}_{2}(\mathbb{R}^{\infty})$ consisting of all 2-planes in $\mathbb{R}^{\infty}$, topologized as the direct limit $\text{Gr}_{2}(\mathbb{R}^{2})\subset\text{Gr}_{2}(\mathbb{R}^{3})\subset\cdots$ (see [MS74] for more details). Identifying each plane with its projection matrix gives an embedding of $\text{Gr}(2)$ into $\mathbb{R}^{\infty\times\infty}$. The total space of the universal plane bundle $\mu$ over $\text{Gr}(2)$ can be modeled by the Stiefel manifold $V(2)=V(2,\infty)$:
$V(2,\infty)=\{A\in\mathbb{R}^{\infty\times 2}:A^{T}A=I\}$
The columns of each $A\in V(2)$ form an orthonormal basis for a plane in $\text{Gr}(2)$ whose projection matrix is given by $AA^{T}$. Thus, the universal plane bundle is then defined by the surjection $\pi_{Gr}:V(2)\to\text{Gr}(2)$ given by $\pi_{Gr}(A)=AA^{T}$.
Given a circle bundle $\xi$ over a paracompact space $B$, the associated plane bundle $\xi_{\text{vect}}$ can be expressed as the pullback of $\mu$ along a classifying map $f:B\to\text{Gr}(2)$. It follows that $\xi$ is isomorphic to the pullback of the associated circle bundle $\mu_{\circ}=V(2)\times_{O(2)}\mathbb{S}^{1}\to\text{Gr}(2)$ along $f$. In this sense $\mu_{\circ}$ is a universal circle bundle, and $\xi$ is classified by the same map as $\xi_{\text{vect}}$.
From the perspective of data science and dimensionality reduction, a major advantage of choosing $\mu_{\circ}$ as our model for the universal circle bundle is that we can apply known techniques to obtain low-dimensional representations of point clouds in $V(2)\times_{O(2)}\mathbb{S}^{1}$. Our coordinatization pipeline for discrete approximate circle bundles integrates Principal Stiefel Coordinates [Lee+25] for this purpose.
@@ 3 Discrete Approximate Circle Bundles
In this section, we introduce discrete approximate notions of local trivializations, Cech cohomology and classifying maps for circle bundles. The purpose of these objects is to represent true circle bundles with finite data which can be algorithmically, efficiently and stably computed from a point cloud and a feature map.
3.1 Three Characterizations
@@
A discrete $(\varepsilon,\delta)$-approximate circle bundle is a map $\pi:X\to B$ between metric spaces together with an open cover $\mathcal{U}=\{U_{j}\}_{j\in J}$ of $B$ and a family of maps $\{f_{j}:\pi^{-1}(U_{j})\to\mathbb{S}^{1}\}_{j\in J}$ with the following properties:
@@
    $d_{H}(f_{l}(\pi^{-1}(U_{jkl})),\mathbb{S}^{1})<\delta$ (with respect to Euclidean distance) for every $U_{jkl}\in\mathcal{N}(\mathcal{U})$.
@@
    whenever $x\in\pi^{-1}(U_{jk})\in\mathcal{N}(\mathcal{U})$.
@@
The family $\{f_{j}\}_{j\in J}$ is called a discrete $(\varepsilon,\delta)$-approximate local trivialization, and the set of discrete $(\varepsilon,\delta)$-approximate local trivializations for $\pi$ subordinate to an open cover $\mathcal{U}$ is denoted by $T_{\varepsilon,\delta}(\pi,\mathcal{U})$. Any $\Omega$ satisfying (10) is said to be a witness that $\{f_{j}\}\in T_{\varepsilon,\delta}(\pi,\mathcal{U})$.
Given a discrete approximate local trivialization, we will refer to the quantity $\alpha\equiv\frac{\varepsilon}{1-\delta}$ as the trivialization quality of the trivialization and write $T_{\alpha}(\pi,\mathcal{U})=\bigcup_{\frac{\varepsilon}{1-\delta}<\alpha}T_{\varepsilon,\delta}(\pi,\mathcal{U})$ for any $\alpha\geq 0$. We define the following metric on $T_{\infty}(\pi,\mathcal{U})$:
$d_{T}(\{f_{j}\}_{j\in J},\{g_{j}\}_{j\in J})=\sup_{j\in J}\sup_{x\in\pi^{-1}(U_{j})}|f_{j}(x)-g_{j}(x)|\\$ (11)
Notice that we require witnesses to take values in $O(2)$ rather than $\text{Homeo}(\mathbb{S}^{1})$.
@@
Next, we introduce discrete and approximate relaxations of classical Cech cohomology. As we have seen, Cech cohomology provides a useful framework for computing obstructions and classifications, so these objects will play a central role in our classification and coordinatization algorithms.
@@
Given a collection of open sets $\mathcal{U}=\{U_{j}\}_{j\in J}$ and a metric group $G$, the set of $G$-valued $\varepsilon$-approximate Cech cocycles subordinate to $\mathcal{U}$ is defined by
$\check{Z}^{k}_{\varepsilon}(\mathcal{U};\mathcal{C}_{G})=\{\Omega\in\check{C}^{k}(\mathcal{U};\mathcal{C}_{G}):\ d_{Z}(\delta^{k}\Omega,\mathbbm{1})<\varepsilon\}$ (12)
where $d_{Z}$ is the supremum metric on $\check{C}^{k}(\mathcal{U};\mathcal{C}_{G})$:
$d_{Z}(\Omega,\Lambda)=\sup_{U_{\sigma}\in\mathcal{N}(\mathcal{U})^{(k)}}\sup_{b\in U_{\sigma}}d_{G}(\Omega_{\sigma}(b),\Lambda_{\sigma}(b))$ (13)
The set of $G$-valued discrete $\varepsilon$-approximate Cech cocycles subordinate to $\mathcal{U}$ is defined by
@@
We will primarily be interested in the case where $G=O(2)$, since every circle bundle over a paracompact space $B$ is represented by a class in $\check{H}^{1}(B;\mathcal{C}_{O(2)})$. Here we endow $O(2)$ with the Frobenius metric, since this is the most natural choice for a number of key results.

### v3

@@ 2.5 The Tautological Circle Bundle Over $\text{Gr}(2)$ And Dimensionality Reduction
$\text{Bun}_{\mathbb{S}^{1}}(B)\quad\leftrightarrow\quad\text{Prin}_{O(2)}(B)\quad\leftrightarrow\quad\text{Vect}_{2,\mathbb{R}}(B)$ (9)
for any paracompact space $B$ (once one fixes the action of $O(2)$ on $\mathbb{S}^{1}$). It follows that up to homotopy, every circle bundle $\xi$ over $B$ can be realized (up to isomorphism) as the sphere bundle of a rank-2 real Euclidean vector bundle $\xi_{\text{vect}}$.
A model for the classifying space of real rank-2 vector bundles is the infinite Grassmann manifold $\text{Gr}(2)=\text{Gr}_{2}(\mathbb{R}^{\infty})$ consisting of all 2-planes in $\mathbb{R}^{\infty}$, topologized as the direct limit $\text{Gr}_{2}(\mathbb{R}^{2})\subset\text{Gr}_{2}(\mathbb{R}^{3})\subset\cdots$ (see [17] for more details). Identifying each plane with its orthogonal projection matrix gives an embedding of $\text{Gr}(2)$ into $\mathbb{R}^{\infty\times\infty}$. The total space of the universal plane bundle $\mu$ over $\text{Gr}(2)$ can be modeled by the Stiefel manifold $V(2)=V(2,\infty)$:
$V(2,\infty)=\{A\in\mathbb{R}^{\infty\times 2}:A^{T}A=I_{2}\}$
The columns of each $A\in V(2)$ form an orthonormal basis for a 2-plane in $\text{Gr}(2)$ whose orthogonal projection matrix is given by $AA^{T}$. Thus, the universal plane bundle is then defined by the surjection
$\begin{array}[]{rccl}\pi_{Gr}:&V(2)&\longrightarrow&\text{Gr}(2)\\ &A&\mapsto&AA^{T}\end{array}$ (10)
Given a circle bundle $\xi$ over a paracompact space $B$, the associated plane bundle $\xi_{\text{vect}}$ (9) can be expressed as the pullback of $\mu$ along a classifying map $f:B\to\text{Gr}(2)$. It follows that $\xi$ is isomorphic to the pullback of the associated circle bundle $\mu_{\circ}=V(2)\times_{O(2)}\mathbb{S}^{1}\to\text{Gr}(2)$ along $f$. Thus, $\mu_{\circ}$ is “universal” in the sense described above, and $\xi$ is classified by the same map as $\xi_{\text{vect}}$.
From the perspective of data science and dimensionality reduction, a major advantage of choosing $\mu_{\circ}$ as our model for the universal circle bundle is that we can apply known dimensionality reduction techniques in $V(2)$ to obtain low-dimensional representations of point clouds in $V(2,d)\times_{O(2)}\mathbb{S}^{1}$, for small $d$. Our coordinatization pipeline for discrete approximate circle bundles (Section 6) integrates Principal Stiefel Coordinates [14] for this purpose.
@@ 3 Discrete Approximate Circle Bundles
In this section, we introduce discrete approximate notions of local trivializations, Čech cohomology and classifying maps for circle bundles. The purpose of these objects is to represent true circle bundles with finite data which can be algorithmically, efficiently and stably computed from a point cloud and a feature map.
3.1 Discrete Approximate Local Trivializations
@@
Let $f:X\to Y$ be a function between metric spaces. The distortion of $f$ is defined by
$\text{dist}(f)=\sup_{x,x^{\prime}\in X}|d_{Y}(f(x),f(x^{\prime}))-d_{X}(x,x^{\prime})|$
Given functions $f:X\to Y$ and $g:Y\to X$, the codistortion of $f$ and $g$ is defined by
$\text{codist}(f,g)=\sup_{x\in X,y\in Y}|d_{Y}(f(x),y)-d_{X}(x,g(y))|$
Definition 3.2.
Let $\pi:X\to B$ and $\pi^{\prime}:X^{\prime}\to B$ be functions between metric spaces. A discrete $(\beta,\varepsilon)$-approximate bundle map between $(X,\pi,B)$ and $(X^{\prime},\pi^{\prime},B)$ is a function $\varphi:X\to X^{\prime}$ with the following properties:
@@
    There exists a function $\psi:X^{\prime}\to X$ such that
    $\max\{\text{dist}(\varphi),\ \text{dist}(\psi),\ \text{codist}(\varphi,\psi)\}\leq\varepsilon,$
@@
    The functions $\varphi$ and $\psi$ satisfy
    $\max\{\sup_{x\in X}d_{B}(\pi(x),\pi^{\prime}\circ\varphi(x)),\quad\sup_{x^{\prime}\in X^{\prime}}d_{B}(\pi^{\prime}(x^{\prime}),\pi\circ\psi(x^{\prime}))\}\leq\frac{\beta}{2}$
Any function $\psi$ satisfying the properties above is called an almost-inverse for $\varphi$. Note that $\psi$ is also a discrete $(\beta,\varepsilon)$-approximate bundle map.
We note several immediate consequences of the definition above for later use:
Proposition 3.3.
Suppose $\varphi:(X,\pi,B)\to(X^{\prime},\pi^{\prime},B)$ is a discrete $(\beta,\varepsilon)$-approximate bundle map and $\psi$ is an almost-inverse for $\varphi$.
  1. 1.
    Almost-Inverse Property:
    $\sup_{x\in X}d_{X}(\psi\circ\varphi(x),x)\leq\varepsilon\hskip 14.22636pt\sup_{x^{\prime}\in X^{\prime}}d_{X}(\varphi\circ\psi(x^{\prime}),x^{\prime})\leq\varepsilon$
  2. 2.
    Restriction: Let $A=\pi^{-1}(U)\subseteq X$ and $A^{\prime}=(\pi^{\prime})^{-1}(U)\subseteq X^{\prime}$ for some set $U\subseteq B$. If $\varphi(A)\subseteq A^{\prime}$ and $\psi(A^{\prime})\subseteq A$, then $\varphi|_{A}$ and $\psi|_{A^{\prime}}$ are almost-inverse discrete $(\beta,\varepsilon)$-approximate bundle maps.
  3. 3.
    Composition: If $\varphi^{\prime}:X^{\prime}\to X^{\prime\prime}$ is a discrete $(\beta^{\prime},\varepsilon^{\prime})$-approximate bundle map from $(X^{\prime},\pi^{\prime},B)$ to $(X^{\prime\prime},\pi^{\prime\prime},B)$ with almost-inverse $\psi^{\prime}$, then $\varphi^{\prime}\circ\varphi$ is a discrete $(\beta+\beta^{\prime},\varepsilon+\varepsilon^{\prime})$-approximate bundle map from $(X,\pi,B)$ to $(X^{\prime\prime},\pi^{\prime\prime},B)$ and $\psi\circ\psi^{\prime}$ is an almost-inverse.
Recall that in the classical setting, a local trivialization for a circle bundle $\pi:X\to B$ is a bundle isomorphism between $(\pi^{-1}(U),\pi|_{\pi^{-1}(U)},U)$ and $(U\times\mathbb{S}^{1},p_{U},U)$ for $U\subseteq B$ open (where $p_{U}$ denotes projection onto the first factor). We therefore turn our attention to discrete approximate bundle maps into product spaces of the form $U\times\mathbb{S}^{1}$.
Definition 3.4.
[…]
More generally, for any $r>0$, let $d_{\mathbb{S}^{1}}^{r}$ denote geodesic distance scaled by a factor of $r$ – that is, $d_{\mathbb{S}^{1}}^{r}(z,z^{\prime})=rd_{\mathbb{S}^{1}}(z,z^{\prime})$ for all $z,z\in\mathbb{S}^{1}$. We will denote the metric space $(\mathbb{S}^{1},d_{\mathbb{S}^{1}}^{r})$ by $\mathbb{S}^{1}_{r}$. Finally, given any metric space $B$, we denote the supremum product metric on $B\times\mathbb{S}^{1}_{r}$ by $d_{\infty}^{r}$.
Remark 3.5.
Since key results related to discrete approximate $O(2)$ Čech cohomology are most easily stated in terms of the Frobenius norm on $O(2)$ (which is closely related to the Euclidean norm on $\mathbb{S}^{1}$), we note the following identity which we will use frequently:
$|z-w|=2\sin^{-1}\left(\frac{d_{\mathbb{S}^{1}}(z,w)}{2}\right),\hskip 28.45274ptz,w\in\mathbb{S}^{1}$ (11)
Definition 3.6.
[…]
Remark 3.7.
[…]
We now introduce the notion of a discrete approximate circle bundle:
Definition 3.8.
A map $\pi:X\to B$ between metric spaces is called a discrete $(r,\beta,\varepsilon)$-approximate circle bundle if there is a good open cover $\mathcal{U}=\{U_{j}\}_{j\in J}$ of $B$ (called a trivializing cover for $(X,\pi,B)$) with the following properties:
  1. 1.
    For each $(jkl)\in\mathcal{N}(\mathcal{U})$, with $j,k,l\in J$ not necessarily distinct, there exists some $x\in X$ such that $\bar{B}_{\beta}(\pi(x))\subseteq U_{jkl}$.
  2. 2.
    For each $j\in J$, there exists a discrete $(r,\beta,\varepsilon)$-approximate local trivialization $\varphi_{j}:\pi^{-1}(U_{j})\to U_{j}\times\mathbb{S}^{1}_{r_{j}}$ for some $r_{j}\geq r$.
The family $\{\varphi_{j}\}_{j\in J}$ is called a discrete $(r,\beta,\varepsilon)$-approximate trivialization of $\pi$, and the set of all such families subordinate to a fixed trivializing cover $\mathcal{U}$ is denoted by $T_{r,\beta,\varepsilon}(\pi,\mathcal{U})$.
Note that any $\Phi=\{\varphi_{j}:\pi^{-1}(U_{j})\to U_{j}\times\mathbb{S}_{r_{j}}^{1}\}_{j\in J}\in T_{r,\beta,\varepsilon}(\pi,\mathcal{U})$ has an associated system of local circular coordinate functions $\mathfrak{F}(\Phi)=\{f_{j}:\pi^{-1}(U_{j})\to\mathbb{S}^{1}\}_{j\in J}$ given by $f_{j}=p_{\mathbb{S}^{1}}\circ\varphi_{j}$. We will see that, as in the classical setting, the relationships between local circular coordinates for a discrete approximate circle bundle encode its global topological structure.
Definition 3.9.
Let $\pi:X\to B$ be a function between metric spaces, and let $\mathcal{U}=\{U_{j}\}_{j\in J}$ be an open cover of $B$. For $\varepsilon>0$ and $0\leq\delta<1$, a system of discrete $(\varepsilon,\delta)$-approximate local circular coordinates for $\pi$ is a family $\mathfrak{F}=\{f_{j}:\pi^{-1}(U_{j})\to\mathbb{S}^{1}\}_{j\in J}$ with the following properties:
  1. 1.
@@
    whenever $x\in\pi^{-1}(U_{jk})$ and $(jk)\in\mathcal{N}(\mathcal{U})$.
  2. 2.
    If $d_{H}$ denotes the Hausdorff distance in $(\mathbb{C},|\cdot|)$, then $d_{H}(f_{l}(\pi^{-1}(U_{jkl})),\mathbb{S}^{1})<\delta$ for every $(jkl)\in\mathcal{N}(\mathcal{U})$.
@@
The set of discrete $(\varepsilon,\delta)$-approximate systems of local angle functions for $\pi$ subordinate to $\mathcal{U}$ is denoted by $\Gamma_{\varepsilon,\delta}(\pi,\mathcal{U})$. Any $\Omega$ satisfying (12) is said to be a witness that $\mathfrak{F}\in\Gamma_{\varepsilon,\delta}(\pi,\mathcal{U})$. Given a system of discrete approximate local circular coordinates, we will refer to the quantity
$\frac{\varepsilon}{1-\delta}$
as the roughness of the system and write
$\Gamma_{\alpha}(\pi,\mathcal{U})=\bigcup\limits_{\frac{\varepsilon}{1-\delta}<\alpha}\Gamma_{\varepsilon,\delta}(\pi,\mathcal{U})$
for any $\alpha>0$. We define the following metric on $\Gamma_{\infty}(\pi,\mathcal{U})$:
$d_{\Gamma}(\{f_{j}\}_{j\in J},\{g_{j}\}_{j\in J})\;\;=\;\;\sup_{j\in J}\sup_{x\in\pi^{-1}(U_{j})}|f_{j}(x)-g_{j}(x)|\\$ (13)
In practice, it will often be more convenient to work with systems of discrete approximate local circular coordinates than discrete approximate local trivializations. Proposition 3.11 below shows that discrete approximate local trivializations always give rise to systems of discrete approximate local circular coordinates. This result is directly implied by the following lemma, which we use again in Section 3.5 to compare systems of local circular coordinates induced by different choices of discrete approximate local trivializations subordinate to the same open cover. The proof is deferred to Appendix A.6.
Lemma 3.10.
Let $\pi:X\to B$ be a function between metric spaces. Suppose $\{\varphi_{j}:(\pi^{-1}(U_{j}),\pi,U_{j})\to(U_{j}\times\mathbb{S}^{1}_{r_{j}},p_{B},U_{j})\}_{j=1}^{2}$ are discrete $(r,\beta,\varepsilon)$-approximate local trivializations for some $r_{1},r_{2}\geq r$ and open sets $U_{1},U_{2}\subseteq B$ such that $\bar{B}_{\beta}(\pi(x))\subseteq U_{12}=U_{1}\cap U_{2}$ for some $x\in X$. Then,
$|r_{1}-r_{2}|\leq\frac{2}{\pi}(\varepsilon+\beta)$
and there exists some $\Omega_{12}\in O(2)$ such that
$|f_{1}(x)-\Omega_{12}f_{2}(x)|\quad\leq\quad\tilde{\tau}(\varphi_{1},\varphi_{2})$
for all $x\in\pi^{-1}(U_{12})$, where $f_{j}=p_{\mathbb{S}^{1}}\circ\varphi_{j}$ for $j=1,2$ and
$\tau(\varphi_{1},\varphi_{2})\ =\ \frac{1}{\min\{r_{1},r_{2}\}}(\text{diam}(U_{12})+10\varepsilon+9\beta),\hskip 28.45274pt\widetilde{\tau}(\varphi_{1},\varphi_{2})\ =\ 2\sin^{-1}\left(\frac{\tau(\varphi_{1},\varphi_{2})}{2}\right)\\$ (14)
The following result is immediate:
Proposition 3.11.
Suppose $\mathcal{U}$ is a trivializing cover for a discrete $(r,\beta,\varepsilon)$-approximate circle bundle $\pi:X\to B$ and $\Phi=\{\varphi_{j}:\pi^{-1}(U_{j})\to U_{j}\times\mathbb{S}^{1}_{r_{j}}\}_{j\in J}\in T_{r,\beta,\varepsilon}(\pi,\mathcal{U})$. The functions $\{f_{j}=p_{\mathbb{S}^{1}}\circ\varphi_{j}\}_{j\in J}$ form a system of discrete $(\tilde{\tau}(\Phi),2\sin^{-1}(\frac{\varepsilon}{2r}))$-approximate local circular coordinates for $\pi$, where
$\tau(\Phi)=\sup_{(jk)\in\mathcal{N}(\mathcal{U})}\tau(\varphi_{j},\varphi_{k}),\hskip 28.45274pt\tilde{\tau}(\Phi)=2\sin^{-1}\left(\frac{\tau(\Phi)}{2}\right)$ (15)
Proof.
Suppose $(jk)\in\mathcal{N}(\mathcal{U})$. If $j<k$, let $\Omega_{jk}\in O(2)$ be any matrix which satisfies
$|f_{j}(x)-\Omega_{jk}f_{k}(x)|\quad\leq\quad\widetilde{\tau}(\varphi_{j},\varphi_{k})\quad\leq\quad\widetilde{\tau}(\Phi)$
for all $x\in\pi^{-1}(U_{jk})$ (such a matrix is guaranteed to exist by Lemma 3.10). If $j>k$, define $\Omega_{jk}=\Omega_{kj}^{-1}$ and observe
$|f_{j}(x)-\Omega_{jk}f_{k}(x)|\quad=\quad|\Omega_{jk}^{-1}f_{j}(x)-f_{k}(x)|\quad=\quad|\Omega_{kj}f_{j}(x)-f_{k}(x)|\quad\leq\quad\widetilde{\tau}(\varphi_{k},\varphi_{j})\quad\leq\quad\widetilde{\tau}(\Phi)$
Lastly, define $\Omega_{jj}=\text{Id}$ for all $j\in J$. Evidently, the matrices $\{\Omega_{jk}\}_{(jk)\in\mathcal{N}(\mathcal{U})}$ can be interpreted as a Čech cochain $\Omega\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$, and we have
$|f_{j}(x)-\Omega_{jk}f_{k}(x)|\ \leq\ \widetilde{\tau}(\Phi)$
whenever $x\in\pi^{-1}(U_{jk})$ for some $(jk)\in\mathcal{N}(\mathcal{U})$.
Now, suppose $U_{jkl}\neq\emptyset$, and let $z\in\mathbb{S}^{1}$ be arbitrary. Let $x\in X$ be so that $\bar{B}_{\beta}(\pi(x))\subseteq U_{jkl}$ and let $b=\pi(x)$. Then,
$d_{\mathbb{S}^{1}}^{r_{j}}(f_{j}(x),z)\quad\leq\quad d_{\infty}^{r_{j}}(\varphi_{j}(x),(b,z))\quad\leq\quad\varepsilon$
so it follows that $d_{H}(f_{j}(\pi^{-1}(U_{jkl})),\mathbb{S}^{1})\leq\frac{\varepsilon}{r_{j}}\leq\frac{\varepsilon}{r}$ with respect to unscaled geodesic distance on $\mathbb{S}^{1}$. Applying identity (11), the Hausdorff bound becomes $2\sin^{-1}(\frac{\varepsilon}{2r})$, so the proposition follows.
Remark 3.12.
In light of the proposition above, we define the roughness of a family $\Phi\in T_{r,\beta,\varepsilon}(\pi,\mathcal{U})$ by
$\alpha(\Phi)=\frac{\widetilde{\tau}(\Phi)}{1-2\sin^{-1}\left(\frac{\varepsilon}{2r}\right)}$ (16)
By the previous result, $\alpha(\Phi)$ is equal to the roughness of $\mathfrak{F}(\Phi)$. Note that our definition of discrete approximate local trivializations requires $r_{j}>\frac{16\varepsilon}{\pi}$ for each $j$, so
$2\sin^{-1}\left(\frac{\varepsilon}{2r}\right)\quad<\quad 2\sin^{-1}\left(\frac{\pi}{32}\right)\quad<\quad\frac{1}{5}$ (17)
for any discrete approximate circle bundle. It follows that $\alpha(\Phi)<\frac{4}{5}\widetilde{\tau}(\Phi)$ for any family of discrete approximate local trivializations.
3.2 Discrete Approximate Čech Cohomology
@@
Here, we introduce discrete and approximate relaxations of classical Čech cohomology. As we have seen, Čech cohomology provides a useful framework for computing obstructions and classifications of circle bundles, so these objects will play a central role in our classification and coordinatization algorithms.
@@
Given a collection $\mathcal{U}=\{U_{j}\}_{j\in J}$ of open subsets of a space $B$ and a metric group $G$, the set of $G$-valued $\varepsilon$-approximate Čech cocycles subordinate to $\mathcal{U}$ is defined by
$\check{Z}^{k}_{\varepsilon}(\mathcal{U};\mathcal{C}_{G})=\left\{\Omega\in\check{C}^{k}(\mathcal{U};\mathcal{C}_{G}):\ d_{Z}(\delta^{k}\Omega,\mathbbm{1})<\varepsilon\right\}$ (18)
where $d_{Z}$ is the supremum metric on $\check{C}^{k+1}(\mathcal{U};\mathcal{C}_{G})$:
$d_{Z}(\Omega,\Lambda)=\sup_{\sigma\in\mathcal{N}(\mathcal{U})^{(k+1)}}\;\sup_{b\in U_{\sigma}}d_{G}(\Omega_{\sigma}(b),\Lambda_{\sigma}(b))$ (19)
and $\mathbbm{1}\in\check{C}^{k+1}(\mathcal{U};\mathcal{C}_{G})$ is the cochain which is constant and equal to the identity of $G$. The set of $G$-valued discrete $\varepsilon$-approximate Čech cocycles subordinate to $\mathcal{U}$ is defined by
@@
We will primarily be interested in the case where $G=O(2)$, since every circle bundle over a paracompact space $B$ is represented by a class in $\check{H}^{1}(B;\mathcal{C}_{O(2)})$. Here we endow $O(2)$ with the Frobenius metric, given its predominance in algorithmic linear algebra and optimization.
3.3 Approximate Classifying Maps